# Probing Reheating Phase via Non-Helical Magnetogenesis and Secondary Gravitational Waves

### v1

@@ II.2.1 Evolution During inflation:
$\rho_{\mathrm{em}}^{\mathrm{I}}=\int_{k_{*}}^{k_{\rm e}}d\ln(k)\times\mathcal{P}_{\rm B/E}(k,\eta_{\rm e})\simeq\frac{H_{{}_{\mathrm{I}}}^{4}}{8\pi}\frac{2^{2|n|+1}\Gamma^{2}\left(|n|+\frac{1}{2}\right)}{\pi^{2}}\frac{\left(1-(k_{*}/k_{\rm e})^{n_{{}_{\mathrm{B}/\mathrm{E}}}}\right)}{n_{{}_{\mathrm{B}/\mathrm{E}}}}.$ (19)
Where $n_{{}_{\mathrm{B}/\mathrm{E}}}=4-2|n|$, represent the spectral indices of the electromagnetic (EM) fields, and $k_{*}$ is the CMB pivot scale. From Eqs. (14)), the spectrum is blue-tilted for $|n|<2$ and red-tilted for $|n|>2$. For scenarios with $|n|\leq 2$, therefore, $\rho_{\rm em}\propto H_{{}_{\mathrm{I}}}^{4}$. In contrast, for $|n|>2$, $\rho_{\rm em}\propto H_{{}_{\mathrm{I}}}^{4}(k_{\rm e}/k_{*})^{|n_{{}_{\mathrm{B}/\mathrm{E}}}|}$, which depends not only on the inflationary energy scale but also on the duration of the inflationary era.
When the energy density of the EM field approaches that of the inflaton, backreaction effects on the background evolution can become significant, affecting expansion dynamics and the power spectrum of primordial perturbations. To quantify this, we define the parameter $\zeta^{\mathrm{I}}=\rho_{\mathrm{em}}^{\mathrm{I}}/\rho_{\mathrm{c}}^{\mathrm{I}}$, providing a measure of the EM field’s influence relative to the background energy. At the end of inflation, the fractional energy density ${\zeta}^{\mathrm{I}}$ is defined as
$\zeta^{\mathrm{I}}\simeq 3.98\times 10^{-12}\left(\frac{H_{{}_{\mathrm{I}}}}{10^{-5}M_{\rm P}}\right)^{2}2^{n+\frac{1}{2}}\Gamma(n+\frac{1}{2})\frac{1-e^{-n_{{}_{\mathrm{B}/\mathrm{E}}}N_{{}_{\mathrm{I}}}}}{n_{{}_{\mathrm{B}/\mathrm{E}}}},$ (20)
When $\zeta^{\mathrm{I}}=1$, the EM energy density becomes comparable to that of the inflationary background, signaling the onset of significant backreaction. For an instantaneous reheating scenario with $N_{{}_{\mathrm{I}}}=55$, with $H_{{}_{\mathrm{I}}}\simeq 10^{-5}M_{\rm P}$, we find that to avoid backreaction, the index parameter must satisfy $|n|\leq 2.18$. This value depends on the inflationary model, as the duration of inflation affects $k_{\rm e}$, which is given by $k_{\rm e}=k_{*}\exp[N_{{}_{\mathrm{I}}}]$.
In the subsequent subsection, we discuss how post-inflationary dynamics, particularly reheating, affect the evolution of primordial magnetic fields. We focus on the role of electrical conductivity during reheating, which plays a crucial role in determining the fate of the generated fields. For simplicity, we consider two extreme scenarios, one with vanishing electrical conductivity and another with very large electrical conductivity. Here we are going to discuss how the zero electrical conductivity limit, the conversion of the electric energy to magnetic energy, can help us to explain the current day observed magnetic field strength at large scale for $n<-1/2$, where the model does not have any kind of backreaction and strong coupling issue.
II.2.2 Evolution During reheating:
Given the assumption of the gauge coupling function $f(\eta)=1$ (see Eq.(11)) after the end of inflation $a>a_{end}$, the evolution equation for the comoving electromagnetic power spectrum during reheating will become
$\mathcal{P}^{c^{\prime}}_{E}=-\mathcal{P}^{c^{\prime}}_{B}$ (21a)
$\mathcal{P}^{c^{\prime\prime}}_{B}+2k^{2}\mathcal{P}^{c}_{B}=2k^{2}\mathcal{P}^{c}_{E}.$ (21b)
@@ II.2.2 Evolution During reheating:
Note that the above equations do not explicitly depend on the scale factor, and that is due to the underlying conformal invariance of the gauge fields. During reheating, there is no coupling term to break the invariance like the one in Eq.(1). So, we can identify the solutions of Eqs(21) as comoving electric and magnetic power spectrum defined as $\mathcal{P}^{c}_{B}=a^{4}\mathcal{P}_{B}$ and $\mathcal{P}^{c}_{E}=a^{4}\mathcal{P}_{E}$. And the solution of the Eqs(21) is governed by
$a^{4}\mathcal{P}_{B}=\frac{d_{2}}{2}+d_{1}\mbox{ cos}(2k(\eta-\eta_{end})),$ (22a)
$a^{4}\mathcal{P}_{E}=\frac{d_{2}}{2}-d_{1}\mbox{ cos}(2k(\eta-\eta_{end}))$ (22b)
where $d_{1},d_{2}$ are the constants, which can be fixed by providing proper initial conditions at the end of inflation, and we get the following expressions,
$d_{1}=\frac{\mathcal{B}-\mathcal{E}}{2}~{}~{};~{}~{}d_{2}=\mathcal{B}+\mathcal{E},$ (23)
where we defined two new constants $\mathcal{B}$ and $\mathcal{E}$ as
$\mathcal{B}=\frac{a_{\mathrm{e}}^{4}\mathcal{P}_{B}^{\mathrm{inf}}(\eta_{\mathrm{end}})}{f^{2}(\eta_{\mathrm{end}})}=\frac{k^{4}}{8\pi}\frac{2^{2n+1}\Gamma^{2}\left(n+\frac{1}{2}\right)}{\pi^{2}}\left(\frac{k}{k_{\rm e}}\right)^{-2n},$ (24)
$\mathcal{E}=\frac{a_{\mathrm{e}}^{4}\mathcal{P}_{E}^{\mathrm{inf}}(\eta_{\mathrm{end}})}{f^{2}(\eta_{\mathrm{end}})}=\frac{k^{4}}{8\pi}\frac{\Gamma^{2}\left(n-\frac{1}{2}\right)2^{2n-1}}{\pi^{2}}\left(\frac{k}{k_{\rm e}}\right)^{2-2n}$ (25)
@@
In de Sitter space, we can identify it through $\left|-\eta_{\mathrm{end}}\right|=1/k_{\rm e}$ where $k_{\rm e}$ is the highest model that crossed the horizon at the end of inflation. The above two Eqs. (24),(25) are evaluated at the end of inflation for the super-horizon modes approximation.

### v2

@@ II.2.2 Backreaction and Strong Coupling Problem
$\rho_{\mathrm{em}}^{\mathrm{I}}=\int_{k_{*}}^{k_{\rm e}}d\ln(k)\times\mathcal{P}_{\rm B/E}(k,\eta_{\rm e})\simeq\frac{H_{{}_{\mathrm{I}}}^{4}}{8\pi}\frac{2^{2|n|+1}\Gamma^{2}\left(|n|+\frac{1}{2}\right)}{\pi^{2}}\frac{\left(1-(k_{*}/k_{\rm e})^{n_{\mathrm{E}}}\right)}{n_{\mathrm{E}}}.$ (21)
To compute the total electromagnetic energy density, we integrate over the comoving wavenumber $k$ in the range $k_{*}\leq k\leq k_{\rm end}$, where $k_{*}$ is the pivot scale corresponding to large wavelengths observable in the CMB, and $k_{\rm end}$ is the smallest scale that exits the horizon at the end of inflation. We take the ultraviolet cutoff to be $k_{\rm UV,c}=k_{\rm end}$, as only those modes that exit the horizon during inflation are excited via the coupling; modes that remain subhorizon throughout inflation stay in the Bunch-Davies vacuum and do not contribute to the observable magnetic field strength.
By requiring that the total energy density of the electromagnetic field remains subdominant compared to the background inflationary energy density at the end of inflation, we find that the coupling parameter must satisfy $n\geq-2.2$, assuming an inflationary energy scale $H_{\rm inf}\simeq 10^{-5}M_{\rm P}$ and an inflationary duration of $N_{\rm inf}\simeq 55$, where $N_{\rm inf}$ denotes the number of e-folds between horizon exit of the pivot scale and the end of inflation. Thus, to simultaneously avoid both the backreaction and strong coupling problems, the allowed range for the coupling parameter is $2.2\leq n<0$. Within this range, the magnetic power spectrum is always blue-tilted. For instance, at n=2n=-2, the electric power spectrum becomes scale-invariant, while the magnetic power spectrum scales as $\mathcal{P}_{B}(k,\eta_{\rm e})\propto k^{2}$, indicating that most of the magnetic energy is concentrated near the cutoff scale $k\simeq k_{\rm end}$. As a result, the amplitude of large-scale magnetic fields is suppressed at the end of inflation.
In the following section, we explore how reheating scenarios involving negligible or low electrical conductivity (EC) can enhance the magnetic field through induction effects (see, for instance, Kobayashi and Sloth (2019)).
II.3 Evolution During reheating phase:
During reheating, we now consider two extreme cases with regard to the electrical conductivity of the background medium.
  1. 1.
    High-Conductivity approximation: This is the conventional and well-studied scenario where all fundamental charged particles are efficiently produced immediately after inflation and get instantaneously thermalised, and the high-temperature background results in a very large electrical conductivity of the universe. In this high-conductivity limit, the rapid response of the charged particles causes the electric field to decay almost instantaneously, effectively removing it from the post-inflationary dynamics. As a result, during reheating, only the magnetic field survives. This can be easily understood from the dynamical equation
    $\mathcal{P}^{c^{\prime}}_{E}+2\sigma\mathcal{P}^{c}_{E}=-\mathcal{P}^{c^{\prime}}_{B}$ (22a)
    $\mathcal{P}^{c^{\prime\prime}}_{B}+\sigma\mathcal{P}^{c^{\prime}}_{B}+2k^{2}\mathcal{P}^{c}_{B}=2k^{2}\mathcal{P}^{c}_{E}.$ (22b)
@@ II.3 Evolution During reheating phase:
    We assume the background conductivity $\sigma$ to be constant. Before we embark, on few points are in order. In the above evolution equation for the comoving electromagnetic power spectrum during reheating, we assume the gauge coupling function $f(\eta)=1$ (see Eq.(11)) after the end of inflation $a>a_{\rm end}$. Further, the equations do not explicitly depend on the scale factor, and that is due to the underlying conformal invariance of the gauge fields. During reheating, there is no coupling term to break the invariance like the one in Eq.(1). Our goal is to work in the large conductivity limit. It can be seen that in the limit $\sigma\rightarrow\infty$, and large scale $k\rightarrow 0$, the consistent solution of the above set of equations would be
    $\mathcal{P}^{c}_{E}(\eta)\approx 0\implies|E|^{2}\approx 0,$
    $\mathcal{P}^{c}_{B}(\eta)\approx\mathcal{B}\implies|B|^{2}\approx\frac{\mathcal{B}}{a(\eta)^{4}}.$
    Where the integration constant $\mathcal{B}=a_{\rm end}^{4}\mathcal{P}^{c}_{B}(\eta_{\rm e})$ is calculated at the time of inflation end. The present-day magnetic field spectrum $\mathcal{P}^{0}_{B}$ with the field strength $B_{0}$ can therefore be calculated as
    $a_{0}^{4}\mathcal{P}^{0}_{B}(k)\simeq a_{\mathrm{re}}^{4}\mathcal{B}\implies B_{0}\simeq\frac{\sqrt{\mathcal{B}}a_{\rm end}^{2}}{{a_{0}^{2}}}.$ (23)
    We restrict our analysis to parameter regimes that avoid both strong coupling and backreaction issues. Since the electric field is no longer present in this case, there is no energy transfer from the electric field to the magnetic field via the Faraday induction mechanism. Therefore, the magnetic field evolves adiabatically for those modes that remain super-horizon during reheating. Therefore, the magnetic energy density redshifts as $a^{-4}$ due to the expansion of the universe. Consequently, the comoving magnetic field power spectrum at the end of reheating is given approximately by $\tilde{\mathcal{P}}_{B}(k,\eta_{\rm reh})\simeq\tilde{\mathcal{P}}_{B}(k,\eta_{\rm e})\,.$ The present-day magnetic field strength at $1\,\mathrm{Mpc}$ scales for a coupling parameter n=2n=-2 is calculated as $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 1.45\times 10^{-36}\,\mathrm{G}\,.$ In contrast, for the lower bound of the coupling parameter, $n\simeq-2.2$, the present-day magnetic field strength is enhanced to $B_{0}(1\,\mathrm{Mpc}^{-1})\simeq 1.3\times 10^{-31}\,\mathrm{G}\,.$ These estimates suggest that for the coupling parameter 2.2<n<0-2.2<n<0, it cannot produce sufficient strength of the magnetic field at $1$ Mpc scale. However, for the case of vanishingly small conductivity, we indeed show that sufficient magnetic field strength can be obtained
  2. 2.
    Low Conductivity approximation: In this case, we assume that the electrical conductivity of the universe remains small during reheating. This assumption is strongly motivated by the fact that, during reheating, the electrically neutral inflaton field is the dominant energy density until the end of reheating. In such a scenario, we are going to discuss how reheating helps us to explain the current observational limit. They have the same evolution equation defined in Eq.(22). As stated earlier, we assume $\sigma\rightarrow 0$ limit. So, we can identify the solutions of Eqs(22) as comoving electric and magnetic power spectrum defined as $\mathcal{P}^{c}_{B}=a^{4}\mathcal{P}_{B}$ and $\mathcal{P}^{c}_{E}=a^{4}\mathcal{P}_{E}$. And the solution of the Eqs(22) is governed by
    $a^{4}\mathcal{P}_{B}=\frac{\mathcal{B}+\mathcal{E}}{2}+\frac{\mathcal{B}-\mathcal{E}}{2}\mbox{ cos}(2k(\eta-\eta_{\rm e})),$ (24a)
    $a^{4}\mathcal{P}_{E}=\frac{\mathcal{B}+\mathcal{E}}{2}-\frac{\mathcal{B}-\mathcal{E}}{2}\mbox{ cos}(2k(\eta-\eta_{\rm e})),$ (24b)
    where $(\mathcal{B},\mathcal{E})$ are the appropriate initial conditions,
    $\mathcal{E}=\frac{a_{\rm end}^{4}\mathcal{P}_{E}^{\mathrm{inf}}(\eta_{\mathrm{end}})}{f^{2}(\eta_{\mathrm{end}})}=\frac{k^{4}}{8\pi}\frac{2^{2n+1}\Gamma^{2}\left(n+\frac{1}{2}\right)}{\pi^{2}}\left(\frac{k}{k_{\rm e}}\right)^{-2n},$ (25)
    $\mathcal{B}=\frac{a_{\rm end}^{4}\mathcal{P}_{B}^{\mathrm{inf}}(\eta_{\mathrm{end}})}{f^{2}(\eta_{\mathrm{end}})}=\frac{k^{4}}{8\pi}\frac{\Gamma^{2}\left(n-\frac{1}{2}\right)2^{2n-1}}{\pi^{2}}\left(\frac{k}{k_{\rm e}}\right)^{2-2n}$ (26)
@@
    In de Sitter space, we can identify it through $\left|-\eta_{\mathrm{end}}\right|=1/k_{\rm e}$ where $k_{\rm end}$ is the highest wavenumber that crossed the horizon at the end of inflation. The above two Eqs. (25),(26) are evaluated at the end of inflation for the super-horizon modes approximation.